# Maria Bazotte, Margarida Carvalho, and Thibaut Vidal

### v3

@@ 5.1.2 Binomial Distribution
${\phi}_{f}^{d}=\frac{1}{\rho}\sum_{i\in\mathcal{N_{F}}}\sum_{p\in\mathcal{P}}\widetilde{\phi}_{fi}^{dp}x_{i}^{p}$ $\forall\ d\in\mathcal{D},\quad f\in\mathcal{N_{F}},$ (19)
where the parameter $\widetilde{\phi}_{fi}^{dp}$ is the impact on the capacity level at facility $f\in\mathcal{N_{F}}$ of installing protection level $p\in\mathcal{P}$ at facility $i\in\mathcal{N_{F}}$ when event $d$ occurs, and $\rho$ is a normalization parameter to guarantee that ${\phi}_{f}^{d}\in[0,1]$. Hence, we adopt convolution for the Binomial distribution in Section 4.4; that is, $\vartheta_{fw}^{s}$ is the realization of a random variable following $U(0,1)$ for scenario $s\in\mathcal{S}$, facility $f\in\mathcal{N_{F}}$, and capacity level $w\in\{1,\ldots,W\}$. The variables $\upsilon^{s}_{fw}\in\{0,1\}$ are equal to 1 if $\vartheta_{fw}^{s}<{\phi}^{\hat{d}(s)}_{f}$ and 0 otherwise for $f\in\mathcal{N_{F}}$ and $w\in\{1,\ldots,W\}$. Constraints (17e) are equivalent to
$\vartheta_{fw}^{s}\upsilon^{s}_{fw}\leq{\phi}_{f}^{\hat{d}(s)}-\epsilon$ $\forall\ f\in\mathcal{N_{F}},\quad w\in\{1,\ldots,W\}$ (20)
$\upsilon^{s}_{fw}\geq{\phi}_{f}^{\hat{d}(s)}-\vartheta_{fw}^{s}$ $\forall\ f\in\mathcal{N_{F}},\quad w\in\{1,\ldots,W\}$ (21)
$\xi_{x}^{fs}=\overline{\nu}\left(\sum_{w=1}^{W}\upsilon_{fw}^{s}\right)$ $\forall\ f\in\mathcal{N_{F}}.$ (22)
Constraints (20) and (21) guarantee the adequate definition of variables $\upsilon^{s}_{fw}$. Constraints (22) determine the capacity value of facilities in every scenario. This variant of the considered problem is denoted NDFPP-Binomial. Note that the original program of this variant is nonlinear nonconvex.
5.1.3 Discrete Distribution
@@
$\hat{p}_{df}^{w}(u)=\frac{u_{df}^{w}}{\sum_{j\in\mathcal{W}}u_{df}^{j}}$ $\ \forall\ d\in\mathcal{D},\quad f\in\mathcal{N_{F}},\quad w\in\mathcal{W},$ (23)
@@
$u_{df}^{w}=\sum_{i\in\mathcal{N_{F}}}\sum_{p\in\mathcal{P}}\widetilde{u}_{fi}^{dpw}x_{i}^{p}$ $\forall\ d\in\mathcal{D},\quad f\in\mathcal{N_{F}},\quad w\in\mathcal{W},$ (24)
where the parameter $\widetilde{u}_{fi}^{dpw}$ is the impact of installing protection $p\in\mathcal{P}$ at facility $i\in\mathcal{N}_{F}$ on the likelihood of capacity level $w\in\mathcal{W}$ at facility $f\in\mathcal{N_{F}}$ when event $d\in\mathcal{D}$ occurs. Thus, we adapt the inverse transformation in Section 4.2.1. For each facility $f\in\mathcal{N_{F}}$ and scenario $s\in\mathcal{S}$, the realization $\vartheta_{f}^{s}$ follows $U(0,1)$. The variables $\theta_{fw}^{s}\in\{0,1\}$ equal one if the capacity level $w\in\mathcal{W}$ is selected for facility $f\in\mathcal{N_{F}}$ in scenario $s\in\mathcal{S}$ and zero otherwise. We define the variables $\pi_{fw}^{s}\in\{0,1\}$, which are equal to $\sum_{j=0}^{w}\theta_{fj}^{s}$, with $\pi_{fW}^{s}$ set to one. This results in the following constraints:
$\pi_{fw}^{s}\sum_{w\in\mathcal{W}}u_{\hat{d}(s)f}^{w}\leq\frac{1}{\vartheta_{f}^{s}}\sum_{j=0}^{w}u_{\hat{d}(s)f}^{j}-\epsilon$ $\forall\ f\in\mathcal{N_{F}},\quad w\in\{0,\ldots,W-1\}$ (25)
$(1-\pi_{f,w-1}^{s})\sum_{w\in\mathcal{W}}u_{\hat{d}(s)f}^{w}\leq\frac{1}{(1-\vartheta_{f}^{s})}\sum_{j=w}^{W}u_{\hat{d}(s)f}^{j}$ $\forall\ f\in\mathcal{N_{F}},\quad w\in\{1,\ldots,W\}$ (26)
$\pi_{fw}^{s}=\pi_{f,w+1}^{s}-\theta_{f,w+1}^{s}$ $\forall\ f\in\mathcal{N_{F}},\quad w\in\{0,\ldots,W-1\}$ (27)
$\sum_{w\in\mathcal{W}}\theta_{fw}^{s}=1$ $\forall\ f\in\mathcal{N_{F}}$ (28)
$\xi_{x}^{fs}=\sum_{w\in\mathcal{W}}\nu_{w}\theta_{fw}^{s}$ $\forall\ f\in\mathcal{N_{F}}.$ (29)
Constraints (25)–(28) are, respectively, the adaptation of Constraints (11)–(14) for the NDFPP. They ensure the correct inverse transformation of $\vartheta_{f}^{s}$ according to the probability distribution defined by Equation (23). Constraints (29) delineate the capacity of facilities according to the capacity level obtained from the proposed transformation. Finally, Constraints (25) and (26) have bilinear terms involving binary variables. To satisfy Assumption 3, we apply the exact McCormick envelope (see Appendix 12 for the linear reformulation). This variant is denoted NDFPP-Discrete. Finally, the original program in this case is also nonlinear nonconvex.
5.1.4 Normal Distribution
@@
$\mu_{f}^{d}=\frac{1}{\rho}\sum_{i\in\mathcal{N_{F}}}\sum_{p\in\mathcal{P}}\widetilde{\mu}_{fi}^{dp}x_{i}^{p}$ $\forall\ d\in\mathcal{D},\quad f\in\mathcal{N_{F}}.$ (30)
@@
The parameter $\widetilde{\mu}_{fi}^{dp}$ is the impact of installing protection $p\in\mathcal{P}$ at facility $i\in\mathcal{N}_{F}$ on the mean of the capacity level of facility $f\in\mathcal{N}_{F}$ when event $d\in\mathcal{D}$ occurs, and $\rho$ is a normalization value. We assume that the capacity is always positive (i.e., $4\widetilde{\sigma}_{f}^{d}\leq\mu_{f}^{d}$ for all feasible decisions). Hence, we use the standardization method in Section 4.3, so the realization $\vartheta_{f}^{s}$ follows $N(0,1)$ for each facility $f\in\mathcal{N_{F}}$ and scenario $s\in\mathcal{S}$. Constraints (17e) are modeled as
$\xi^{fs}_{x}=\mu_{f}^{\hat{d}(s)}+\vartheta_{f}^{s}\sigma_{f}^{\hat{d}(s)}$ $\forall\ f\in\mathcal{N_{F}}.$ (31)
This problem variant is denoted NDFPP-Normal, and its original program is nonlinear nonconvex.
6 Experimental Results
@@
This section validates the application of our methodology for addressing 2-SMIPD. First, Section 6.1 summarizes the instance generation. Section 6.2 then compares the transformed program for NDFPP-Selection solved with the SAA and the DEP of its original program, specifically, the MILP reformulation obtained by auxiliary linearizations in Bhuiyan et al. (2020). The DEP of the original program for the NDFPP-Binomial, NDFPP-Discrete, and NDFPP-Normal is nonlinear nonconvex (see Section 5). Thus, Section 6.3 only explores the performance of the estimators obtained by the SAA applied to our transformed program. Finally, we investigate the value of the stochastic solution (VSS) of the NDFPP in Section 6.4. Note that different methods can be used to solve our obtained two-stage program with exogenous uncertainty. We apply the SAA given its simplicity and desirable properties (see Section 3.1).
@@
The instances of the SAA program, the DEP, the EV problem, and the expected result of the EV solution (EEV) for computing the VSS were coded in C++ and solved directly with Gurobi 10.0.3. All experiments were processed on the Beluga cluster of the Compute Canada network on an Intel Gold 6148 Skylake @ 2.4 GHz with a maximum of 100 GB. We limited Gurobi to using one thread. The details on the application of the SAA are given in Appendix 8. The code is available on GitHub (https://github.com/mariabazotte/saa_endog).
6.1 Instance Generation
We follow a procedure similar to that of Bhuiyan et al. (2020) for network generation. We use the latitude, longitude, and population data of cities in the southeastern United States from the basic US Cities dataset, which is available at https://simplemaps.com/data/us-cities. First, we select cities with a population exceeding a predefined threshold as the nodes of the network. Then, departing from Bhuiyan et al. (2020), we use Delaunay triangulation to define the edges of the network. Thus, we do not restrict the edge lengths and the number of neighbors for nodes. We investigate five networks, with $|\mathcal{N}_{\mathcal{T}}^{0}|=15$, 25, 30, 39, and 48. The thresholds and number of edges for these networks are presented in Appendix 15. The transportation cost $q_{ij}$ of arcs $(i,j)\in\mathcal{A}\backslash\mathcal{A}_{m}$ is the geometric distance $\varrho_{e(i,j)}$ between cities $i$ and $j$. The demands are set to the node population divided by $10^{4}$. We set the dummy facility cost multiplier $a$ to $10$. Following Bhuiyan et al. (2020), we solve the capacitated facility location and network design problem from Melkote and Daskin (2001) to define the set of facilities. For this, we use the median home value from www.zillow.com, 2023, and we set the node capacity to ${\sum_{n\in\mathcal{N}_{\mathcal{T}}^{0}}b_{n}}/{0.9|\mathcal{N}_{\mathcal{T}}^{0}|}$. We consider the values 4 and 5 for $|\mathcal{N}_{\mathcal{F}}|$. This code is available from the same GitHub project.
@@ 6.1 Instance Generation
Analogously to Bhuiyan et al. (2020), we define the probability of success of increasing the capacity level by one $\overline{\phi}_{f}^{dp}$ as $({0.95\times c_{f}^{p}}/{c_{f}^{P}})^{{\hat{l}(f,d)}/{L}}$, where $\hat{l}(f,d)$ is the intensity level in $\mathcal{L}$ at which facility $f\in\mathcal{N_{F}}$ is impacted by disruption $d\in\mathcal{D}_{0}$. However, we consider a 5% risk of protection malfunction on the success of one capacity level. For the no-disruption ($nd$) event, we assign $\overline{\phi}_{f}^{nd,p}$ the value $\eta({0.95\times c_{f}^{p}}/{c_{f}^{P}})+(1-\eta)$, where $\eta=0.7$ represents the degree to which the protection decisions impact the uncertainty in the capacity level for this event. Bhuiyan et al. (2020) consider that facilities are not impacted by the no-disruption ($nd$) event and by disruption events occurring in low-intensity-level regions. Thus, we have instances with a larger number of scenarios.
For the NDFPP-Binomial and NDFPP-Discrete, we define $\widetilde{\phi}_{fi}^{dp}$ and $\widetilde{u}_{fi}^{dpw}$ as $\overline{\phi}_{i}^{dp}$ if $f=i$, and as $0.3\cdot\overline{\phi}_{i}^{dp}$ otherwise. For the NDFPP-Binomial and NDFPP-Normal, we define $\rho=1+0.3\cdot(|\mathcal{N_{F}}|-1)$. For the NDFPP-Normal, we set $\widetilde{\mu}_{fi}^{dp}$ to $\overline{\mu}_{f}^{dp}$ if $f=i$, and to $0.3\cdot\overline{\mu}_{f}^{dp}$ otherwise. The parameter $\overline{\mu}_{f}^{dp}$ is defined as $\nu_{\text{max}}\left[1-\frac{(P+1-p)[L+\hat{l}(f,d)+2]}{2(P+1)(L+1)}\right]$ for each protection level $p\in\mathcal{P}$, where $\hat{l}(f,d)$ returns zero for the no-disruption event ($nd$). Thus, the mean increases with increasing protection and decreases with increasing intensity levels. The standard deviation $\widetilde{\sigma}^{d}_{f}$ is given by $\frac{1}{\rho}\left[\overline{\sigma}_{f}^{d}+0.3\cdot\sum_{i\in\mathcal{N}_{F}\backslash\{f\}}\overline{\sigma}_{i}^{d}\right]$ where $\overline{\sigma}_{f}^{d}$ is defined as $\nu_{\text{max}}\left[\frac{L+\hat{l}(f,d)+2}{8(P+1)(L+1)}\right]$. Hence, it increases with increasing intensity levels and is independent of protection. This way, $0\leq 4\overline{\sigma}_{f}^{d}\leq\overline{\mu}_{f}^{d1}\leq\cdots\leq\overline{\mu}_{f}^{dP}$.
6.2 Comparative results for NDFPP-Selection
The DEP of the original program is formulated in Bhuiyan et al. (2020) as an MILP obtained by linearizing the probability function (4). To solve this MILP, we implemented the same methodology as Bhuiyan et al. (2020): an L-shaped decomposition accelerated with nondominated cuts and valid inequalities. We compare this approach with solving our transformed program with the SAA. Note that we cannot apply the SAA to the original program as it has endogenous uncertainty.

### v4

@@ 5.1 Network Design and Facility Protection Problem
${\phi}_{f}^{d}=\frac{1}{\rho}\sum_{i\in\mathcal{N_{F}}}\sum_{p\in\mathcal{P}}\widetilde{\phi}_{fi}^{dp}x_{i}^{p}\ \quad\forall\ d\in\mathcal{D},\quad f\in\mathcal{N_{F}},$ (17)
where the parameter $\widetilde{\phi}_{fi}^{dp}$ is the impact on the capacity level at facility $f\in\mathcal{N_{F}}$ of installing protection level $p\in\mathcal{P}$ at facility $i\in\mathcal{N_{F}}$ when event $d$ occurs, and $\rho$ is a normalization parameter to guarantee that ${\phi}_{f}^{d}\in[0,1]$. Hence, we employ the convolution technique for the Binomial distribution in Section 4.4; that is, $\vartheta_{fw}^{s}$ is the realization following the $U(0,1)$ for scenario $s\in\mathcal{S}$, facility $f\in\mathcal{N_{F}}$, and capacity level $w\in\{1,\ldots,W\}$. The variables $\upsilon^{s}_{fw}\in\{0,1\}$ are equal to one if $\vartheta_{fw}^{s}<{\phi}^{\hat{d}(s)}_{f}$ and zero otherwise for $f\in\mathcal{N_{F}}$ and $w\in\{1,\ldots,W\}$. Finally, we adapt Constraints (14) to guarantee the adequate definition of variables $\upsilon^{s}_{fw}$ (see Appendix 12). Thus, we determine the capacity value in every scenario as $\xi_{x}^{fs}=\overline{\nu}\left(\sum_{w=1}^{W}\upsilon_{fw}^{s}\right)$ for every facility $f\in\mathcal{N_{F}}$, which are equivalent to Constraints (16e). This variant of the considered problem is denoted NDFPP-Binomial. Note that the DE-OP of this variant is nonlinear nonconvex (see Appendix 13).
5.1.3 Discrete Distribution.
@@
$\hat{p}_{df}^{w}(u)=\frac{u_{df}^{w}}{\sum_{j\in\mathcal{W}}u_{df}^{j}}\ \quad\forall\ d\in\mathcal{D},\quad f\in\mathcal{N_{F}},\quad w\in\mathcal{W},$ (18)
@@
$u_{df}^{w}=\sum_{i\in\mathcal{N_{F}}}\sum_{p\in\mathcal{P}}\widetilde{u}_{fi}^{dpw}x_{i}^{p}\ \quad\forall\ d\in\mathcal{D},\quad f\in\mathcal{N_{F}},\quad w\in\mathcal{W},$ (19)
where the parameter $\widetilde{u}_{fi}^{dpw}$ is the impact of installing protection $p\in\mathcal{P}$ at facility $i\in\mathcal{N}_{\mathcal{F}}$ on the likelihood of capacity level $w\in\mathcal{W}$ at facility $f\in\mathcal{N_{F}}$ when event $d\in\mathcal{D}$ occurs. Thus, we adapt the inverse transformation in Section 4.1.1. For each facility $f\in\mathcal{N_{F}}$ and scenario $s\in\mathcal{S}$, the realization $\vartheta_{f}^{s}$ follows the $U(0,1)$. The variables $\theta_{fw}^{s}\in\{0,1\}$ equal one if the capacity level $w\in\mathcal{W}$ is selected for facility $f\in\mathcal{N_{F}}$ in scenario $s\in\mathcal{S}$ and zero otherwise. We also define the variables $\pi_{fw}^{s}\in\{0,1\}$, which are equal to $\sum_{j=0}^{w}\theta_{fj}^{s}$, with $\pi_{fW}^{s}$ set to one. We adapt the Constraints (10)–(13) for the NDFPP to ensure the correct inverse transformation of $\vartheta_{f}^{s}$ according to the probability distribution defined by Equation (18). The obtained constraints initially contain bilinear terms involving binary variables, and thus, to satisfy Assumption 3, we apply the exact McCormick envelope. This adaptation and linearization procedures are detailed in Appendix 12. As a result, we have $\xi_{x}^{fs}=\sum_{w\in\mathcal{W}}\nu_{w}\theta_{fw}^{s}$ for every facility $f\in\mathcal{N_{F}}$, which model Constraints (16e). This variant is denoted NDFPP-Discrete. Finally, the DE-OP in this case is also nonlinear nonconvex (see Appendix 13).
5.1.4 Normal Distribution.
@@
$\mu_{f}^{d}=\frac{1}{\rho}\sum_{i\in\mathcal{N_{F}}}\sum_{p\in\mathcal{P}}\widetilde{\mu}_{fi}^{dp}x_{i}^{p}\ \quad\forall\ d\in\mathcal{D},\quad f\in\mathcal{N_{F}}.$ (20)
@@
The parameter $\widetilde{\mu}_{fi}^{dp}$ is the impact of installing protection $p\in\mathcal{P}$ at facility $i\in\mathcal{N}_{\mathcal{F}}$ on the mean of the capacity level of facility $f\in\mathcal{N}_{\mathcal{F}}$ when event $d\in\mathcal{D}$ occurs, and $\rho$ is a normalization value. We ensure that the capacity is always positive, i.e., $\mu_{f}^{d}\geq 4\widetilde{\sigma}_{f}^{d}$ for all feasible decisions. We use the standardization method in Section 4.3, so the realization $\vartheta_{f}^{s}$ follows $N(0,1)$ for each facility $f\in\mathcal{N_{F}}$ and scenario $s\in\mathcal{S}$. Constraints (16e) are modeled as $\xi^{fs}_{x}=\mu_{f}^{\hat{d}(s)}+\vartheta_{f}^{s}\sigma_{f}^{\hat{d}(s)}$ for every $f\in\mathcal{N_{F}}$. This variant is denoted NDFPP-Normal, and its DE-OP is nonlinear nonconvex (see Appendix 13).
5.2 Experimental Results
@@
This section validates the application of our methodology for addressing 2-SMIPD. First, Section 5.2.1 summarizes the instance generation. Section 5.2.2 then compares the SAA applied to the transformed program (SAA-TP) and the DE-OP for NDFPP-Selection, specifically, the MILP reformulation obtained by auxiliary linearizations in Bhuiyan et al. (2020). The DE-OP for the NDFPP-Binomial, NDFPP-Discrete, and NDFPP-Normal is nonlinear nonconvex (see Section 5). Thus, Section 5.2.3 only explores the performance of the estimators obtained by the SAA-TP. Finally, we investigate the value of the stochastic solution (VSS) of the NDFPP in Section 5.2.4. Note that different methods can be used to solve our obtained two-stage program with exogenous uncertainty. We apply the SAA, given its simplicity and desirable properties (see Section 3.1).
@@
The instances of the transformed SAA program, the DE-OP, the EV problem, and the expected result of the EV solution (EEV) for computing the VSS were coded in C++ and solved directly with Gurobi 10.0.3. All experiments were processed on the Beluga cluster of the Compute Canada network on an Intel Gold 6148 Skylake @ 2.4 GHz with a maximum of 100 GB. We limited Gurobi to using one thread. The details on the application of the SAA are given in Appendix 7. All instances, results, and code are publicly available on GitHub (https://github.com/mariabazotte/saa_endog).
5.2.1 Instance Generation.
We follow a procedure similar to that of Bhuiyan et al. (2020) for network, transportation costs, and client demand generation, which is detailed in Appendix 16 and available on the same GitHub project. We consider five networks, with $|\mathcal{N}_{\mathcal{T}}^{0}|=15$, 25, 30, 39, and 48, and assume the values 4 and 5 for $|\mathcal{N}_{\mathcal{F}}|$. The protection costs $c_{f}^{p}$, which are randomly generated, and the edge-opening costs $c_{e}$ are defined similarly as in Bhuiyan et al. (2020) (see Appendix 16). The budget $C$ is a percentage of the total cost with maximum protection, $C=0.5(\sum_{n\in\mathcal{N_{F}}}c_{n}^{P}+\sum_{e\in\mathcal{E}}c_{e})$, making it adjustable to different instance sizes. We consider four protection levels ($P=4$), and define the maximum facility capacity as $\nu_{\text{max}}=|{\sum_{n\in\mathcal{N_{C}}}b_{n}}/{0.9|\mathcal{N_{F}}|}|$. Thus, the capacity increase $\overline{\nu}$ is ${\nu_{\text{max}}}/{W}$ with capacity values $\nu_{w}=w\overline{\nu}$ for $w\in\mathcal{W}$. We use $W=2,3,$ and $4$. As in Bhuiyan et al. (2020), we consider three disruption events with distinct centers of occurrence ($|\mathcal{D}_{0}|=3$), and we assume three disruption intensity levels ($L=3$). We use the procedure in Bhuiyan et al. (2020) to define the intensity level $\hat{l}(f,d)$ in $\mathcal{L}$ at which facility $f\in\mathcal{N_{F}}$ is impacted by disruption $d\in\mathcal{D}_{0}$ (see Appendix 16).
Analogously to Bhuiyan et al. (2020), we define the probability of successfully increasing the capacity level by one, $\overline{\phi}_{f}^{dp}$, as $({0.95\times c_{f}^{p}}/{c_{f}^{P}})^{{\hat{l}(f,d)}/{L}}$, with a 5% risk of protection malfunction for each capacity level increase. For the $nd$ event, $\overline{\phi}_{f}^{nd,p}$ is defined as $\eta({0.95\times c_{f}^{p}}/{c_{f}^{P}})+(1-\eta)$, where $\eta=0.7$ reflects the influence of protection decisions on capacity level uncertainty on this event. Unlike Bhuiyan et al. (2020), we assume that facilities are affected by the $nd$ event and low-intensity disruption events, which results in instances with a larger number of scenarios. For the NDFPP-Binomial and NDFPP-Discrete, we set $\widetilde{\phi}_{fi}^{dp}$ and $\widetilde{u}_{fi}^{dpw}$ to $\overline{\phi}_{i}^{dp}$ if $f=i$, and to $0.3\times\overline{\phi}_{i}^{dp}$ otherwise. For the NDFPP-Binomial and NDFPP-Normal, $\rho$ is defined as $1+0.3\times(|\mathcal{N_{F}}|-1)$. For the NDFPP-Normal, $\widetilde{\mu}_{fi}^{dp}$ is $\overline{\mu}_{f}^{dp}$ if $f=i$, and $0.3\times\overline{\mu}_{f}^{dp}$ otherwise. The parameter $\overline{\mu}_{f}^{dp}$ is given by $\nu_{\text{max}}\left[1-\frac{(P+1-p)(L+\hat{l}(f,d)+2)}{2(P+1)(L+1)}\right]$ for each protection level $p\in\mathcal{P}$, with $\hat{l}(f,d)$ returning zero for the $nd$ event, meaning that the mean increases with protection and decreases with intensity. The standard deviation $\widetilde{\sigma}^{d}_{f}$ is $\frac{1}{\rho}\left[\overline{\sigma}_{f}^{d}+0.3\sum_{i\in\mathcal{N}_{\mathcal{F}}\backslash\{f\}}\overline{\sigma}_{i}^{d}\right]$, where $\overline{\sigma}_{f}^{d}$ is $\nu_{\text{max}}\left[\frac{L+\hat{l}(f,d)+2}{8(P+1)(L+1)}\right]$. This standard deviation increases with intensity levels and is independent of protection, ensuring $4\overline{\sigma}_{f}^{d}\leq\overline{\mu}_{f}^{d1}\leq\cdots\leq\overline{\mu}_{f}^{dP}$.
5.2.2 Comparative results for NDFPP-Selection.
For the case of discrete selection of distributions, the literature on endogenous uncertainty has mainly linearized the probability function (3) with different approaches. Here, we consider the DE-OP formulated in Bhuiyan et al. (2020) as an MILP. To solve this MILP, we implemented the same methodology as Bhuiyan et al. (2020): an L-shaped decomposition accelerated with nondominated cuts and valid inequalities. We compare this approach with the SAA-TP. Note that we cannot apply the SAA to the DE-OP as it has endogenous uncertainty.
@@
0010101010202020203030303040404040505050506060606070707070808080809090909010010010010000101010102020202030303030404040405050505060606060707070708080808090909090100100100100Instances (%)SAA-TP $\left(\overline{gap}(\overline{x})(\%)\right)$DE-OP $\left(GAP(\%)\right)$
Comparison of the optimality gap obtained by the SAA-TP and the DE-OP.
Table 5.2.2 shows the experimental results of the 4- and 5-facility instances. In the table, $|K|$ and $|K_{t}|$ are the set of scenarios in the DE-OP (Program (2)) and in the deterministic equivalent of our transformed Program (4), respectively. Moreover, $GAP(\%)$ is the optimality gap for the DE-OP. For the DE-OP, all runs reached the time limit of two hours. We followed the procedure explained in Section 3.1 for the SAA-TP to obtain the feasible solution $\overline{x}$ and its estimators (see Appendix 7 for details). Even though the number of scenarios $|K_{t}|$ of the transformed program is extremely large (i.e.,$|\mathcal{D}|{|\mathcal{W}|}^{|\mathcal{N_{F}}||\mathcal{P}|}$), the number $|K|$ of scenarios of the DE-OP is small (i.e., $|\mathcal{D}|{|\mathcal{W}|}^{|\mathcal{N_{F}}|}$) for the considered instances. Thus, we can compute the real stochastic value of the feasible solution $\overline{x}$, i.e., $v(\overline{x})=c^{\top}\overline{x}+g(\overline{x})$. For the SAA-TP, we solve $M=50$ problems with $N=750$ sample scenarios because these values sufficed to obtain estimators with low variance and feasible, high-quality solutions. We present the over-estimator of the relative optimality gap, i.e., $\overline{gap}(\overline{x})(\%)={100\%\times[v(\overline{x})-\overline{v}_{M}^{N}]}/{v(\overline{x})}$. Because the lower-bound estimator $\overline{v}_{M}^{N}$ is a random variable, this over-estimator of the optimality gap may assume negative values. The computational time of the SAA-TP is in seconds. For all runs, the SAA-TP finished before the time limit of two hours. For comparison, we display the gap between the upper-bound values obtained by the SAA-TP and the DE-OP, i.e., $GAP_{UB}(\%)={100\%\times[UB-v(\overline{x})]}/{UB}$, and the gap between their lower bounds, i.e., $GAP_{LB}(\%)={100\%\times[LB-\overline{v}_{M}^{N}]}/{LB}$, where $UB$ and $LB$ are the upper and lower bounds for the DE-OP, respectively. Next, we summarize our main findings.
Result 1
The SAA-TP produces a feasible solution $\overline{x}$ with a considerably low over-estimator of the optimality gap, $\overline{gap}(\overline{x})(\%)$ is between $0\%$ and $1\%$ for all instances. The lower bound $\overline{v}_{M}^{N}$ and the optimality gap estimators have low variability, as shown by the standard deviation $\hat{\sigma}_{N,M}$ of the lower-bound estimator, which is inferior to $0.5\%$ of $\overline{v}_{M}^{N}$ for all instances (see Appendix 17 for details). For the considered instances, the computational time required by the SAA-TP was inferior to 2000s in 87% of the cases, with an average of around 990s.
Result 2
The SAA-TP’s over-estimator of the optimality gap is significantly smaller than the DE-OP’s optimality gap (the difference is around $24\%$ on average). The SAA-TP obtained these optimality gaps in significantly less time than the DE-OP, and, within the time limit, consistently found feasible solutions with slightly better values than those from the DE-OP.
Result 1 highlights the good performance of the SAA-TP. Result 2 shows that it outperformed the DE-OP. Figure 5.2.2, which shows the optimality gap for the DE-OP and the SAA-TP, corroborates this analysis. The SAA-TP obtained optimality gaps close to $0\%$ for all instances, whereas the DE-OP obtained optimality gaps inferior to $10\%$ only for $22\%$ of the instances. The DE-OP obtained optimality gaps of $10\%$$20\%$ for around $56\%$ of the instances, of $20\%$$30\%$ for around $8\%$ of the instances, and of $100\%$ for around $14\%$ of the instances. Even though the transformed Program (4) has more scenarios than the DE-OP, the SAA-TP solves different problem instances with fewer scenarios. Therefore, although solving the DE-OP is constrained to a relatively small number of scenarios, the transformed program can handle larger instances as we can use sampling techniques to address it. The SAA-TP obtained high-quality solutions by sampling only $N=750$ scenarios for each of the $M=50$ SAA problems, which, in most cases, is less than the number $|K|$ of scenarios of the DE-OP. Despite the number of scenarios, the SAA-TP solves mathematical programs smaller than the DE-OP (in terms of number of variables and constraints). The DE-OP’s MILP formulation contains numerous big-M constraints due to the linearization of the probability function (3), which leads to weak linear relaxations. In contrast, the transformation applied to obtain the NDFPP-Selection formulation (see Section 5.1.1) includes no extra variables or constraints (see Appendix 18 for a comparison of the number of variables and constraints in the DE-OP and the SAA-TP). Finally, although the SAA performed well in this case study, other sampling methods can be employed based on the problem’s specific requirements.
5.2.3 Analysis for NDFPP-Binomial, NDFPP-Discrete, and NDFPP-Normal.
In this section, we analyze the statistical estimators and feasible solution $\overline{x}$ of the SAA-TP for NDFPP-Binomial, NDFPP-Discrete, and NDFPP-Normal, which are MILPs (see Sections 5.1.2-5.1.4). In these cases, the DE-OP (Program (2) for discrete distributions) is nonlinear nonconvex, and the linearization of Bhuiyan et al. (2020) cannot be applied. Moreover, note that our transformed Program (4) for these variants has continuous random variables, making the explicit enumeration of scenarios impossible and requiring the implementation of sampling techniques. The DE-OP has a small number $|K|$ of scenarios for the NDFPP-Binomial and NDFPP-Discrete instances. Thus, we compute $v(\overline{x})$ for these cases. The number of scenarios of the DE-OP is infinite for NDFPP-Normal because the endogenous random variables are continuous. Consequently, we compute the estimator $\hat{v}_{N{{}^{\prime}}}(\overline{x})$ of $v(\overline{x})$. Tables 5.2.3 and 5.2.3 lists the experimental results of the 4- and 5-facility instances. The over-estimator of the optimality gap $\overline{gap}(\overline{x})(\%)$ is defined as in the previous section for NDFPP-Binomial and NDFPP-Discrete. For NDFPP-Normal, $\overline{gap}(\overline{x})(\%)={100\%\times[\hat{v}_{N{{}^{\prime}}}(\overline{x})-\overline{v}_{M}^{N}]}/{\hat{v}_{N{{}^{\prime}}}(\overline{x})}$. We set $M=50$ and $N=750$ for the same reason as in the previous section. For NDFPP-Normal, we use $N{{}^{\prime}}=5\times 10^{4}$. The SAA-TP finished for all instances before reaching 3 hours. The result is as follows:
Result 3
The SAA-TP obtained a feasible solution $\overline{x}$ with a significantly low statistical over-estimator of the optimality gap $\overline{gap}(\overline{x})(\%)$ between $-1\%$ and $1\%$ for all instances of the problem versions. The SAA’s estimators have high confidence, as shown by the standard deviation $\hat{\sigma}_{N,M}$ of the lower-bound estimator, which is inferior to $0.5\%$ of $\overline{v}_{M}^{N}$ for all cases, and the standard deviation $\hat{\sigma}_{N{{}^{\prime}}}(\overline{x})$ of the upper bound estimator, which is inferior to $0.2\%$ of $\hat{v}_{N{{}^{\prime}}}(\overline{x})$ for all NDFPP-Normal instances (see Appendix 19 for details). The computational time was, on average, around 1500s for NDFPP-Binomial, 3000s for NDFPP-Discrete, and 1200s for NDFPP-Normal.
[…]
Experimental results for NDFPP-Binomial and NDFPP-Discrete of the SAA-TP.